ut -kuxx = F(x, t) 0<x<l t>0

u(x, 0)= f(x) for 0<x<l and u(0, t)= u(l, t)= 0 for t >0


1
Expert's answer
2021-02-04T08:07:37-0500

CORRECTED SOLUTION

We seek the solution of the form of the functions that are solutions for the homogenous equation:

"u(t,x)=\sum\limits_{n=1}^{+\infty}u_n(t)\sin\frac{\pi n}{l}x"

Then

"u_t = \sum\limits_{n=1}^{+\infty}u'_n(t)\sin\frac{\pi n}{l}x"

"u_{xx}=-\sum\limits_{n=1}^{+\infty}(\frac{\pi n}{l})^2 u_n(t)\sin\frac{\pi n}{l}x"

"u_t-ku_{xx} = \sum\limits_{n=1}^{+\infty}(u'_n(t)+k(\frac{\pi n}{l})^2 u_n(t))\sin\frac{\pi n}{l}x"

Let "F(t,x)=\sum\limits_{n=1}^{+\infty}F_n(t)\sin\frac{\pi n}{l}x" (a Fourier series for the function F(t,x) on the interval "0\leq x\leq l" )

then "f_n(t)=\frac{2}{l}\int\limits_0^l F(t,x)\sin\frac{\pi n}{l}x\, dx"

Substituting this to the equation "u_t-ku_{xx}=F(t,x)" we have:

"\sum\limits_{n=1}^{+\infty}(u'_n(t)+k(\frac{\pi n}{l})^2 u_n(t)-F_n(t))\sin\frac{\pi n}{l}x=0"

At the left side there is a Fourier series for the zero function, hence, all the Fourier coefficients are zeros, that is,

"u'_n(t)+k(\frac{\pi n}{l})^2u_n(t)=F_n(t)"

The solution of the homogeneous equation (with "F_n(t)=0" ) is

"u_n(t)=c_ne^{-k(\pi n/l)^2t}"

To solve the non-homogeneous equation we are varying the constant cn .

"u_n(t)=c_n(t)e^{-k(\pi n/l)^2t}"

We have the equation (ODE) for cn(t):

"c'_n(t)e^{-k(\pi n/l)^2t} = F_n(t)"

To find the initial conditions we consider the function

"f(x) = u(0,x) =\sum\limits_{n=1}^{+\infty}u_n(0)\sin\frac{\pi n}{l}x=\sum\limits_{n=1}^{+\infty}c_n(0)\sin\frac{\pi n}{l}x"

Therefore, "c_n(0)=\frac{2}{l}\int\limits_0^l f(x)\sin\frac{\pi n}{l}x\, dx" - Fourier coefficients for the function "f(x)" .

Integrating ODE we have

"c_n(t)=c_n(0)+\int\limits_0^t F_n(s)e^{ks(\pi n/l)^2}ds"


Answer. "u(t,x)=\sum\limits_{n=1}^{+\infty}c_n(t)e^{-k(\pi n/l)^2t}\sin\frac{\pi n}{l}x"

with

"c_n(t)=c_n(0)+\int\limits_0^t F_n(s)e^{ks(\pi n/l)^2}ds"

"F_n(t)=\frac{2}{l}\int\limits_0^l F(t,x)\sin\frac{\pi n}{l}x\, dx"

"c_n(0)=\frac{2}{l}\int\limits_0^l f(x)\sin\frac{\pi n}{l}x\, dx"